OpenStudy (sleepyjess):

Help with a quick derivative problem? Find the derivative of f(x) = \(\dfrac 5x\) at x = -1.

OpenStudy (anonymous):

\[f(x)=\frac{ 5 }{x}=5x ^{-1}\]

OpenStudy (anonymous):

Now you can use the power rule

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=5(x)^{-1} }\) \(\large\color{black}{ \displaystyle f'(x)=\color{red}{(-1)\times }5(x)^{-1\color{red}{-1}} }\)

OpenStudy (solomonzelman):

simplify that...

OpenStudy (sleepyjess):

I'm not sure if I'm familiar with the power rule

OpenStudy (solomonzelman):

oh, I can introduce you then:) ((( I am using a d/dx notation to denote a derivative, or a notation of f'(x) for a derivative of some function f(x) ))) \(\large\color{black}{ \displaystyle \frac{\rm d }{ {\rm d}x} \left(x^{\rm n}\right)={\rm n}\cdot x^{{\rm n}-1} }\) (this is the power rule)

OpenStudy (solomonzelman):

the above rule applies to any power n, besides n=0.

OpenStudy (sleepyjess):

Okay, that doesn't seem too bad :)

OpenStudy (solomonzelman):

yes, not bad at all. can you write the power rule for x^3 ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{\rm d }{ {\rm d}x} \left(x^{\rm 3}\right)=~? }\) what would that be?

OpenStudy (sleepyjess):

\(3*x^2\)?

OpenStudy (solomonzelman):

yes, \(3x^2\)

OpenStudy (sleepyjess):

Okay, not too bad at all :) Seem pretty easy right now actually :3

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{\rm d }{ {\rm d}x} \left(x^{\rm 3}\right)={\rm 3} x^{2} }\) (that line says: "the derivative of x^3 is equal to 3x^2" )

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

now, go ahead and do the power rule for \(x^{-1}\): \(\large\color{black}{ \displaystyle \frac{\rm d }{ {\rm d}x} \left(x^{\rm -1}\right) }\)

OpenStudy (kainui):

Just to make sure, are you expected to take the derivative by the limit definition? That would explain why you don't know the power rule.

OpenStudy (solomonzelman):

oh... I haven't thought that way because i read "a `quick` derivative problem" :)

OpenStudy (sleepyjess):

It doesn't tell us a specific way to do it, just wants us to find the answer somehow

OpenStudy (solomonzelman):

ok, so then using power rule would be simpler

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{\rm d }{ {\rm d}x} \left(x^{\rm \color{red}{-1}}\right)={\rm \color{red}{(-1)}}\cdot x^{{\rm \color{red}{-1}}-1} }\)

OpenStudy (solomonzelman):

now, for any coefficient c in front of \(x^{\rm n}\), (of course c\(\ne\)0) the derivative is: \(\large\color{black}{ \displaystyle \frac{\rm d }{ {\rm d}x} \left({\rm c}\cdot x^{\rm n}\right)={\rm c}\cdot {\rm n}\cdot x^{{\rm n}-1} }\)

OpenStudy (solomonzelman):

in your case the power of x - the n, is -1. & the coefficient c in front is 5.

OpenStudy (sleepyjess):

\(-5x^{-2}\)

OpenStudy (solomonzelman):

yup

OpenStudy (solomonzelman):

or, \(\large\color{black}{ \displaystyle \frac{ -5 }{ x^2} }\)

OpenStudy (solomonzelman):

now, you need to plug in -1 for x.

OpenStudy (sleepyjess):

so, 5?

OpenStudy (solomonzelman):

no exactly:

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{-5}{x^2} }\) is the derivative. plug in -1 for x \(\large\color{black}{ \displaystyle \frac{-5}{(-1)^2} }\) and you get ?

OpenStudy (sleepyjess):

-5/1, so 5!

OpenStudy (sleepyjess):

*-5

OpenStudy (solomonzelman):

yes, -5 (but don't put the exclamation mark, because then it is a factorial of 5 times -1) jk

OpenStudy (solomonzelman):

f`(-1)=-5

OpenStudy (solomonzelman):

Saying , that the slope of the function f(x)=5/x, at x=-1 is equivalent to -5.